OpenStudy (anonymous):

Find the slope of the line through the pair of points. A(2, –3), P(2, 9)

OpenStudy (anonymous):

i just wanna know if the answer is 0

OpenStudy (jlastino):

y2-y1/x2-x1 (9+3)/(2-2) 12/0 undefined. It forms a vertical line

OpenStudy (anonymous):

slope of a line is given by : y2-y1/x2-x1 =9+3/2-2=12/0 and as Jlastino said , its not defined because 0 in the denominator means that its infinity, which in turn is not defined